Given that cos x = 15/17 and sin x < 0, then csc x =
Answer by lwsshak3(11628)

Given that cos x = 15/17 and sin x < 0, then csc x =
***
Reference angle in quadrant IV where cos>0, sin<0.

csc(x)=1/sin(x)=-17/8